# LLM-SAP: Large Language Model Situational Awareness Based Planning

### v3

@@ 4.3 LLM Evaluators
Experiments are conducted to assess the feasibility of using LLMs as evaluators to score generated FSM plans. As shown in Table 4, the ability of GPT-4 and Claude-2 to rank FSM outputs similarly to human rankings is tested.
The models are tasked with ranking FSM plans in pairs, groups of 4, and groups of 6. Rankings are then compared to expert human rank orders to quantify accuracy. Tests find both GPT-4 and Claude-2 could rank FSM pairs with 75.7% agreement to human ranking, evidencing reliability for comparative evaluation. However, accuracy drops significantly for ranking 4 or more FSMs. Table 5 and Fig. 4 show that GPT-4 and Claude-2 aligned best with human judgment when evaluating outputs from the models themselves. For example, GPT-4’s scoring of GPT-4 FSMs closely matches expert ranking. This demonstrates LLMs can provide sound comparative assessments, particularly for their own model family’s outputs. The experiments illustrate LLMs’ potential to serve as evaluators that mimic human appraisals of planning formalisms. By relying on relative rather than absolute assessments, variability is reduced.
  RBS Scores Human GPT-4 Claude-2
GPT-4+SAP 1.13 1.15 1.42
GPT-4 1.88 1.85 1.58
GPT-3.5+SAP 1.25 1.35 1.04
GPT-3.5 1.75 1.65 1.96
Claude-2+SAP 1.14 1.30 1.23
Claude-2 1.86 1.70 1.77
Accuracy 100% 75.70% 75.70%
  RBS Scores Claude2 Human
GPT-3.5+SAP+feedback 1.38 1.35
GPT-4+SAP 1.62 1.65
GPT-3.5+SAP+feedback 1.00 1.00
GPT-3.5+SAP 2.00 2.00
@@ 4.4 Multi-Agent Improvement
A closed-loop experiment is conducted to quantify improvements in planning through iterative generation and evaluation between two agents. GPT-3.5 with the SAP prompt is utilized as the generative model (LLM_gen) and Claude-2 served as the evaluative model (LLM_eval).
LLM_gen first proposes an FSM, which LLM_eval then scores and provides feedback. Incorporating this feedback, LLM_gen produced an enhanced FSM in the next round. Testing shows the updated FSM design exceeded the initial round’s quality with a higher RBS score after just one iteration. As shown in Table 6, the closed-loop FSM surpasses even GPT-4 with the SAP prompt, previously the top standalone performer. Further analysis reveals the feedback-improved output contains more planning details, leading to increased RBS scores. This demonstrates two weaker models can effectively boost each other’s deficiencies through collaborative dialogue. The results empirically prove interactive cycles of generative prototyping and critical evaluation between LLMs can unlock latent strengths. By leveraging complementary capacities, the multi-agent approach strengthens deductive reasoning and planning beyond individual model limits.
@@ 5 Conclusion
Our analysis provides initial evidence that situational awareness prompts and multi-agent feedback improve planning reliability and humaneness. Experiments on a home hazard benchmark show gains when prompting relationship reasoning and consequence consideration, as exemplified by the planning outputs in Appendix C. Results specifically demonstrate that GPT-3.5, GPT-4, and Claude-2 have superior deduction and mapping with open-ended prompts, especially for complex safety scenarios. Additionally, a complete closed-loop experiment proves that multi-agent collaboration boosts quality. While advances are promising, efficiently condensing fluid cognition absent environmental feedback remains an open challenge. Important future work includes expanded planning corpus training and coupling deduction with simulation. As capabilities progress, introduced techniques appear useful for instilling robustness. However, oversight guardrails remain vital for ensuring ethical alignment before real-world instantiation. By assessing generative planning capacities and limitations, this research steers progress at the intersection of deduction, awareness, and embodiment while balancing innovation with ethical caution.
@@ A.2 Human Annotator Evaluation
As shown in Table 8,9 and 10, the table is divided into three parts, mainly displaying the details of the rankings of each question measured by humans and the mean of the overall rankings. For more specific content, such as the best demonstration written by humans and detailed descriptions of scenes, please check our GitHub for detailed content. The annotation ”None” indicates that the model cannot output at this time due to its limitations.
@@ A.3 Planning Qualitative Analysis
A review of the state machines produced by large language models on this dataset reveals several common shortcomings compared to human-authored finite state machines. GPT-4 demonstrates reasonably coherent state transitions and actions in constrained situations but still falls short of human planning, especially long-term motivations. GPT-4 can understand scenarios comparably to GPT-4+SAP, with logical reasoning, but lacks care-oriented foresight, risking harm. GPT-3.5+SAP and GPT-3.5 show tendencies for unclear situational comprehension, chaotic logic, and state machine instability compared to GPT-4. GPT-3.5 compliance suffers regarding action specificity, partially mitigated by SAP. While Claude-2+SAP attempts simple state changes, safety awareness and self-limitations are often lacking, improved somewhat with SAP but still inconsistent. Claude-2 displays straightforward but limited state transitions, with unconventional logic and unresolved scenarios upon analysis.
@@ Appendix B Ablation Studies
B.1 Ablation Study of One-shot Impact
@@ C.1 Example 1 in scene 10
Fig. 8: Scene description: In this home study, a clear mobility challenge is presented: a person in a wheelchair is positioned at the entrance, facing an office chair that has been left carelessly in the middle of the room, blocking the path to the desk. The wheelchair user, ready to work or study, appears momentarily halted by this obstacle. The study itself is neatly arranged with a desk against the far wall, flanked by bookshelves filled with books and a filing cabinet to the side. The desk is set up for use, with a computer and some papers neatly placed on it. However, the thoughtlessly placed office chair in the centre of the room creates a barrier, preventing direct access from the doorway to the desk. The room is brightly lit, with sunlight streaming through a window and illuminating the space, casting a shadow of the wheelchair onto the floor. This lighting also highlights the obstructive office chair, underscoring the ease with which it could be moved.
@@ C.2 Example 2 in scene 18
Fig. 11: Scene description: In this living room scene, a young child is playing on the floor close to an uncovered electrical outlet. The child, curious and unaware of the danger, is reaching towards the outlet with a metal toy in hand, posing a significant risk of electric shock. The room is comfortably furnished with a couch, a colourful rug, and various toys spread around, indicating that it’s a space commonly used for play. However, the presence of the accessible, unprotected outlet in the play area makes it a safety concern for the child. Bright, natural light fills the room from a nearby window, drawing attention to the child’s innocent exploration and the looming danger of the electrical outlet.
@@ C.3 Example 3 in scene 6
Fig. 14: Scene description: This scene is set in a well-lit laundry room. On the floor, a variety of household cleaning products are carelessly left open and within easy reach. There’s a bucket with a mop, suggesting recent cleaning activity, but the chlorine bleach bottle next to it is uncapped, and a detergent container is tipped over, its contents slowly oozing out. Shelves above the washing machine hold more chemicals, like fabric softeners and stain removers, with some of their caps loosely placed or missing. The laundry room’s door is ajar, suggesting easy access, especially for curious pets or children. The stark lighting in the room highlights the vivid labels on the cleaning products, many of which feature bright, eye-catching designs, potentially attracting unintended interest.
@@ C.4 Example 4 in scene 20 from closed-loop multi-agent
Fig. 17: Scene description: In this living room scene, an elderly person is seated in a recliner, appearing to be in distress and in need of assistance. Across the room, an emergency response system device is placed on a high shelf, well out of reach. The person is trying to get up from the chair, reaching out towards the device, but it’s clear they are unable to stand and walk across the room to access it. The room is comfortably furnished with a couch, a coffee table, and family photos on the walls, creating a cozy atmosphere. However, the placement of the emergency device in a hard-to-reach location poses a significant risk, as the elderly individual cannot easily activate it in case of an emergency. Natural light filters in through a window, casting a gentle glow in the room and highlighting the elderly person’s struggle and the poorly placed emergency device.

### v5

@@ 4.3 LLM Evaluators
Experiments assess the feasibility of using LLMs, like GPT-4 and Claude-2, to score FSM plans and compare their rankings with human evaluations, as shown in Appendix Table 9. The models are tested with ranking FSM plans in pairs, as well as in groups of 4 and 6. These rankings are then compared to expert human rank orders to measure accuracy. Tests find both GPT-4 and Claude-2 could rank FSM pairs with 75.7% agreement to human ranking, evidencing reliability for comparative evaluation. However, their accuracy significantly decreases when ranking groups of 4 or more FSMs. Table 10 and Fig. 6 in Appendix B show that GPT-4 and Claude-2 align most closely with human judgment when evaluating outputs generated by the models themselves. For example, GPT-4’s scoring of its own FSMs closely matches expert rankings. This demonstrates that LLMs can provide accurate comparative assessments, particularly for outputs from their own model family. The experiments highlight the potential of LLMs to serve as evaluators that mimic human appraisals of planning formalisms. By focusing on relative rather than absolute assessments, variability is minimized.
@@ 4.4 Multi-Agent Improvement
A closed-loop experiment quantifies planning improvements through iterative generation and evaluation between two agents. GPT-3.5 with the SAP prompt serves as the generative model (LLM_gen), and Claude-2 serves as the evaluative model (LLM_eval). LLM_gen initially proposes an FSM, which LLM_eval scores and provides feedback on. Using this feedback, LLM_gen produces an improved FSM in the next round. Testing shows the updated FSM surpasses the initial quality, achieving a higher RBS score after one iteration. Appendix Table 11 shows the closed-loop FSM outperforms GPT-4 with the SAP prompt, the previous top standalone performer. The feedback-improved output features more detailed planning, boosting RBS scores. This demonstrates how two weaker models can compensate for each other’s shortcomings through collaboration. The results indicate interactive cycles between LLMs enhance reasoning and planning by leveraging their complementary strengths, surpassing individual model capabilities.
@@ 5 Conclusion
This study marks an advancement in the field of situational awareness-based planning using LLMs. New benchmarks, a specialized dataset, and multi-agent strategies have improved the planning capabilities of LLMs, better equipping them to handle complex and unpredictable human-centric scenarios. Looking ahead, further explorations will focus on expanding datasets and refining model architectures to speed up reasoning, with a particular emphasis on evaluating end-to-end VLMs to bridge the time-lag gap between simulated and real-time environments. This research underscores the importance of stimulating latent reasoning in LLMs and paves the way for ethically sound and safe AI planning processes in practical applications.
@@ A.2 Human Annotator Evaluation
Levels Quantity Description
Level 1 9 Interacting with objects alone.
Level 2 4 Simple human interactions with predictable objects.
Level 3 5 Hard to predict people and animals, indirect actions.
Level 4 6 Highly urgent scenarios.
Total 24 24 scenarios as the evaluation set.
As shown in Table 2,3 and 4, the table is divided into three parts, mainly displaying the details of the rankings of each question measured by humans and the mean of the overall rankings. For more specific content, such as the best demonstration written by humans and detailed descriptions of scenes, please check our GitHub for detailed content. The annotation ”None” indicates that the model cannot output at this time due to its limitations.
@@ A.3 Planning Qualitative Analysis
A review of the state machines produced by LLMs on this dataset reveals several common shortcomings compared to human-authored FSMs. GPT-4 demonstrates reasonably coherent state transitions and actions in constrained situations but still falls short of human planning, especially long-term motivations. GPT-4 can understand scenarios comparably to GPT-4+SAP, with logical reasoning, but lacks care-oriented foresight, risking harm. GPT-3.5+SAP and GPT-3.5 show tendencies for unclear situational comprehension, chaotic logic, and state machine instability compared to GPT-4. GPT-3.5 compliance suffers regarding action specificity, partially mitigated by SAP. While Claude-2+SAP attempts simple state changes, safety awareness and self-limitations are often lacking, improved somewhat with SAP but still inconsistent. Claude-2 displays straightforward but limited state transitions, with unconventional logic and unresolved scenarios upon analysis.
@@ Appendix B Ablation Studies
  State Coverage
Transition Coverage
Cyclomatic Complexity
Safety Mindset
Scalability
Assistance UX
Actions Set Alignment
@@
B.1 Ablation Study of One-shot Impact
  RBS Scores Human annotated Rankings
Overall level 1 level 2 level 3 level 4
GPT-4+SAP 1.21 1.22 1.00 1.00 1.50
GPT-4 2.04 1.78 2.00 2.40 2.17
GPT-3.5+SAP 3.08 3.11 3.25 3.40 2.67
GPT-3.5 3.71 4.00 3.75 3.20 3.67
Claude-2+SAP 5.09 5.00 5.25 5.00 5.17
Claude-2 5.86 5.86 5.75 6.00 5.83
  Ranking Tasks Ranking Accuracy
GPT-4 Claude-2
Pairs
53/70 (75.7%)
53/70 (75.7%)
Group of 4
4/46 (8.6%)
4/46 (8.6%)
Group of 6
0/22 (0%)
1/22 (4.5%)
  RBS Scores Human GPT-4 Claude-2
GPT-4+SAP 1.13 1.15 1.42
GPT-4 1.88 1.85 1.58
GPT-3.5+SAP 1.25 1.35 1.04
GPT-3.5 1.75 1.65 1.96
Claude-2+SAP 1.14 1.30 1.23
Claude-2 1.86 1.70 1.77
Accuracy 100% 75.70% 75.70%
  RBS Scores Claude2 Human
GPT-3.5+SAP+feedback 1.38 1.35
GPT-4+SAP 1.62 1.65
GPT-3.5+SAP+feedback 1.00 1.00
GPT-3.5+SAP 2.00 2.00
@@ C.1 Example 1 in scene 10
Fig. 8: Scene description: In this home study, a clear mobility challenge is presented: a person in a wheelchair is positioned at the entrance, facing an office chair that has been left carelessly in the middle of the room, blocking the path to the desk. The wheelchair user, ready to work or study, appears momentarily halted by this obstacle.
The study itself is neatly arranged with a desk against the far wall, flanked by bookshelves filled with books and a filing cabinet to the side. The desk is set up for use, with a computer and some papers neatly placed on it. However, the thoughtlessly placed office chair in the centre of the room creates a barrier, preventing direct access from the doorway to the desk.
The room is brightly lit, with sunlight streaming through a window and illuminating the space, casting a shadow of the wheelchair onto the floor. This lighting also highlights the obstructive office chair, underscoring the ease with which it could be moved.
@@ C.2 Example 2 in scene 18
Fig. 11: Scene description: In this living room scene, a young child is playing on the floor close to an uncovered electrical outlet. The child, curious and unaware of the danger, is reaching towards the outlet with a metal toy in hand, posing a significant risk of electric shock.
The room is comfortably furnished with a couch, a colourful rug, and various toys spread around, indicating that it’s a space commonly used for play. However, the presence of the accessible, unprotected outlet in the play area makes it a safety concern for the child.
Bright, natural light fills the room from a nearby window, drawing attention to the child’s innocent exploration and the looming danger of the electrical outlet.
@@ C.3 Example 3 in scene 6
Fig. 14: Scene description: This scene is set in a well-lit laundry room. On the floor, a variety of household cleaning products are carelessly left open and within easy reach. There’s a bucket with a mop, suggesting recent cleaning activity, but the chlorine bleach bottle next to it is uncapped, and a detergent container is tipped over, its contents slowly oozing out.
Shelves above the washing machine hold more chemicals, like fabric softeners and stain removers, with some of their caps loosely placed or missing. The laundry room’s door is ajar, suggesting easy access, especially for curious pets or children.
The stark lighting in the room highlights the vivid labels on the cleaning products, many of which feature bright, eye-catching designs, potentially attracting unintended interest.
@@ C.4 Example 4 in scene 20 from closed-loop multi-agent
Fig. 17: Scene description: In this living room scene, an elderly person is seated in a recliner, appearing to be in distress and in need of assistance. Across the room, an emergency response system device is placed on a high shelf, well out of reach.
The person is trying to get up from the chair, reaching out towards the device, but it’s clear they are unable to stand and walk across the room to access it. The room is comfortably furnished with a couch, a coffee table, and family photos on the walls, creating a cosy atmosphere. However, the placement of the emergency device in a hard-to-reach location poses a significant risk, as the elderly individual cannot easily activate it in case of an emergency.
Natural light filters in through a window, casting a gentle glow in the room and highlighting the elderly person’s struggle and the poorly placed emergency device.